OpenStudy (nuccioreggie):

Variable p is 2 more than variable d. Variable p is also 1 less than variable d. Which pair of equations best models the relationship between p and d? p = d + 2 p = d − 1 p = d − 2 p = d + 1 d = 2p d = 2p − 1 d = 2p d = 2p + 1

OpenStudy (nuccioreggie):

@welshfella

OpenStudy (nuccioreggie):

@jameshorton

OpenStudy (jameshorton):

yea

OpenStudy (jameshorton):

Variable p is 2 more than variable d. Variable p means p is means equals "=" 2 more than means add 2 or plus 2 p = 2 + d

OpenStudy (jameshorton):

does that make sense?

OpenStudy (nuccioreggie):

yea so A

OpenStudy (jameshorton):

for the second part "Variable p is 1 less than variable d." Variable p is means "p =" We want one less than variable d, that means d -1 Combining this we have: p = d - 1

OpenStudy (nuccioreggie):

so A

OpenStudy (jameshorton):

yep

OpenStudy (nuccioreggie):

i have few more question can you help me wit those

OpenStudy (jameshorton):

sure